# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 2.3 Linear perturbations
Following Ref. HVWp1 , we take small perturbations on top of the rigidly-rotating state. For definiteness, we consider only longitudinal perturbations along the vorticity vector, such that $\delta u^{\mu}=\delta u\delta^{\mu}_{z}$. The detailed derivation of the equations for the linear perturbations is presented in Ref. HVWp1 , and for the sake of brevity, we do not repeat it here. Instead, we list the results:
@@
where $D=u^{\mu}\partial_{\mu}$ represents the comoving derivative, $\nabla^{\mu}=\Delta^{\mu\nu}\partial_{\nu}$ is the spatial gradient in the fluid rest frame, while $\theta=\partial_{\mu}u^{\mu}$ is the expansion scalar. As we already discussed in Sec. 2.2 at the level of the charge currents, all three currents $J^{\mu}_{V/A/H}$ would be conserved in a quantum field theory of free (non-interacting) massless fermions. This is certainly true also in the quantum case for the vector current when $\partial_{\mu}J^{\mu}_{V}=0$. Contrary to Ref. HVWp1 , in this paper, we take into account the non-conservation of the axial and helical charges in the so-called relaxation-time approximation.
@@
where $\omega(k)$ represent the angular frequencies determined by the system of Eqs. (10). [^5] It is important not to confuse the angular frequency $\omega=\omega(k)$, which determines the time evolution of fluctuations (11), with the vorticity $\omega^{\mu}$ which enters, for example, the system of equations (10e) being associated with the local angular velocity of the vortical fluid.
In order for linear perturbations to be applicable, we consider an unpolarized background state in which $\mu_{A}=\mu_{H}=0$. Then, the right-hand sides of Eqs. (10d) and (10e) are also linear in perturbations. Overall, considering the Fourier mode characterized by $k$ and $\omega(k)$, Eqs. (10a) and (10b) reduce to
$-3\omega\delta P_{\omega}+4kP_{\omega}\delta u_{\omega}=0,\qquad k\delta P_{\omega}-4\omega P_{\omega}\delta u_{\omega}=0,$ (12a)
while the equations for the charge currents become
$\sum_{\ell^{\prime}=V,A,H}\left(\omega\delta_{\ell,\ell^{\prime}}-k\Omega\frac{\partial\sigma^{\omega}_{\ell}}{\partial Q_{\ell^{\prime}}}\right)\delta Q_{\ell^{\prime}}-k\Omega\frac{\partial\sigma^{\omega}_{\ell}}{\partial P}\delta P-(kQ_{\ell}-2\omega\Omega\sigma_{\ell}^{\omega})\delta u$ $=-\frac{iT^{2}\delta\mu_{\ell}}{3\tau_{\ell}},$ (12b)
where we set $\tau_{V}\rightarrow\infty$ as the vector charge is conserved.
The perturbations in the energy-momentum sector, described by Eq. (12a), are not influenced by perturbations in the charges, $\delta Q_{\ell}$. The non-trivial solutions of Eqs. (12a) correspond to the acoustic modes,
@@
with $c_{s}$ being the speed of sound in an ultrarelativistic ideal fluid. For the charge modes, Eqs. (12a) enforce $\delta P_{\omega}=\delta u_{\omega}=0$, such that Eq. (12b) becomes:
$\bigl{(}\omega\delta Q_{\ell}-k\Omega\delta\sigma^{\omega}_{\ell}\bigr{)}{\biggl{|}}_{\delta P=0}=-\frac{iT^{2}\delta\mu_{\ell}}{3\tau_{\ell}}.$ (14)
The above equation involves the variations $\delta Q_{\ell}$ and $\delta\sigma^{\omega}_{\ell}$ of the charge densities and vortical conductivities at constant pressure. Thinking in terms of the grand canonical ensemble parameters, the constraint $\delta P=0$, together with the thermodynamic relation $\delta P=s\delta T+\sum_{\ell}Q_{\ell}\delta\mu_{\ell}$ leads to the constraint
$\delta T=-\sum_{\ell}\frac{Q_{\ell}}{s}\delta\mu_{\ell}.$ (15)
@@
$\mathbb{M}_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime}}=0,\qquad\mathbb{M}_{\ell\ell^{\prime}}=\omega T^{2}\mathbb{M}^{\omega}_{\ell\ell^{\prime}}-k\Omega T\mathbb{M}^{\Omega}_{\ell\ell^{\prime}}+\frac{iT^{2}}{3\tau_{\ell}}\delta_{\ell\ell^{\prime}},$ (16a)
where we introduced the following notations:
$\mathbb{M}^{\omega}_{\ell\ell^{\prime}}$ $=\frac{1}{T^{2}}\left(\frac{\partial Q_{\ell}}{\partial\mu_{\ell^{\prime}}}-\frac{3Q_{\ell}Q_{\ell^{\prime}}}{sT}+\frac{Q_{\ell^{\prime}}\vec{\mu}}{sT}\cdot\frac{\partial Q_{\ell}}{\partial\vec{\mu}}\right),$ (16b)
$\mathbb{M}^{\Omega}_{\ell\ell^{\prime}}$ $=\frac{1}{T}\left(\frac{\partial\sigma^{\omega}_{\ell}}{\partial\mu_{\ell^{\prime}}}-\frac{2\sigma^{\omega}_{\ell}Q_{\ell^{\prime}}}{sT}+\frac{Q_{\ell^{\prime}}\vec{\mu}}{sT}\cdot\frac{\partial\sigma^{\omega}_{\ell}}{\partial\vec{\mu}}\right).$ (16c)
@@ 2.4 Unpolarized background
The pressure corresponding to an ensemble of non-interacting, massless fermions and anti-fermions with polarizations $\lambda=\pm 1/2$ is given by Ambrus:2019khr ; HVWp1 :
@@
$q^{V}_{\sigma,\lambda}=\sigma,\qquad q^{A}_{\sigma\lambda}=2\lambda,\qquad q^{H}_{\sigma,\lambda}=2\lambda\sigma.$ (19)
@@
$\sigma^{\omega}_{\ell}$ $=\frac{1}{2}\frac{\partial Q_{\ell}}{\partial\mu_{A}}=-\frac{T^{2}}{2\pi^{2}}\sum_{\sigma,\lambda}2\lambda q_{\sigma,\lambda}^{\ell}{\rm Li}_{2}(-e^{\mu_{\sigma,\lambda}/T}).$ (21)
@@
${\rm Li}_{2}(-e^{a})+{\rm Li}_{2}(-e^{-a})$ $=-\frac{\pi^{2}}{6}-\frac{a^{2}}{2},$ (24)
@@
$P$ $=\frac{7\pi^{2}T^{4}}{180}+\frac{\mu_{V}^{2}T^{2}}{6}+\frac{\mu_{V}^{4}}{12\pi^{2}},$ $Q_{V}$ $=\frac{\mu_{V}T^{2}}{3}+\frac{\mu_{V}^{3}}{3\pi^{2}},$
$\sigma_{H}^{\omega}$ $=\frac{T^{2}}{\pi^{2}}\left[{\rm Li}_{2}(-e^{-\mu_{V}/T})-{\rm Li}_{2}(-e^{\mu_{V}/T})\right],$ $\sigma_{A}^{\omega}$ $=\frac{T^{2}}{6}+\frac{\mu_{V}^{2}}{2\pi^{2}},$ (25)
@@
We now divide Eq. (16) by $T^{2}$ and obtain
$\left[\omega\mathbb{M}^{\omega}_{\ell\ell^{\prime}}-\kappa_{\Omega}\mathbb{M}^{\Omega}_{\ell\ell^{\prime}}+\frac{i}{3\tau_{\ell}}\delta_{\ell,\ell^{\prime}}\right]\delta\mu_{\ell^{\prime}}=0,$ (27)
where we introduced the parameter
$\kappa_{\Omega}=\frac{k\Omega}{T}\,,$ (28)
which has a sense of a normalized momentum. Taking into account the following expressions for the derivatives of the charge densities and vortical conductivities with respect to the chemical potentials,
$\frac{\partial Q_{\ell}}{\partial\mu_{\ell^{\prime}}}{\biggl{|}}_{\mu_{A}=\mu_{H}=0}$ $=2\begin{pmatrix}\sigma^{\omega}_{A}&0&0\\ 0&\sigma^{\omega}_{A}&\sigma^{\omega}_{H}\\ 0&\sigma^{\omega}_{H}&\sigma^{\omega}_{A}\end{pmatrix},$ $\frac{\partial\sigma^{\omega}_{\ell;\beta}}{\partial\mu_{\ell^{\prime}}}$ $=\frac{1}{\pi^{2}}\begin{pmatrix}0&\mu_{V}&TL\\ \mu_{V}&0&0\\ TL&0&0\end{pmatrix},$ (29)
with $L=2\ln\left(2\cosh\frac{\mu_{V}}{2T}\right)$, we find for $\mathbb{M}_{\omega}$ and $\mathbb{M}_{\Omega}$ the following expressions HVWp1 :
$\mathbb{M}_{\omega}$ $=\frac{2}{T^{2}}\begin{pmatrix}\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H&0&0\\ 0&\sigma^{\omega}_{A}&\sigma^{\omega}_{H}\\ 0&\sigma^{\omega}_{H}&\sigma^{\omega}_{A}\end{pmatrix},$ $\mathbb{M}_{\Omega}$ $=\begin{pmatrix}0&\frac{1}{H}A&\frac{1}{H}B\\ A&0&0\\ B&0&0\end{pmatrix}.$ (30)
In the above, $\Delta H=H-1$ with $H=(e+P)/sT=1+\mu_{V}Q_{V}/(sT)$, while $A$ and $B$ are defined as
@@
$\sigma_{A}^{\omega}=\frac{T^{2}}{6}+\frac{\mu_{V}^{2}}{2\pi^{2}},\qquad\sigma^{\omega}_{H}=\frac{2\ln 2}{\pi^{2}}T^{2}\alpha_{V}+O(\alpha_{V}^{3}),\qquad H=1+\frac{15\alpha_{V}^{2}}{7\pi^{2}}+O(\alpha_{V}^{4}),$
$L=2\ln 2+O(\alpha_{V}^{2}),\quad A=\frac{2\alpha_{V}}{7\pi^{2}}+O(\alpha_{V}^{3}),\quad B=\frac{2\ln 2}{\pi^{2}}+\frac{7\pi^{2}-120\ln 2}{28\pi^{4}}\alpha_{V}^{2}+O(\alpha_{V}^{4}).$ (32)
At large chemical potential, when $|\alpha_{V}|\rightarrow\infty$, we have
$H=\frac{\alpha_{V}^{2}}{\pi^{2}}\left(1+\frac{23\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right),\qquad A=\frac{2\alpha_{V}}{3\pi^{2}}\left(1-\frac{4\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right),$
$\frac{4s_{V}Q_{v}}{\pi^{2}s}=\frac{4s_{V}}{\pi^{2}\alpha_{V}}(H-1)\simeq\frac{4s_{V}\alpha_{V}}{\pi^{4}}\left(1+\frac{8\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right).$ (33a)
In what concerns the quantities $L$, $B$ and $\sigma_{H}$, their large-$\alpha_{V}$ behaviour receives exponentially-damped corrections of the form $e^{-|\alpha_{V}|}$, which can be accessed taking into account the sign $s_{V}={\rm sgn}(\mu_{V})$ of $\mu_{V}$. Then, one obtains trivially:
$L\simeq|\alpha_{V}|+2e^{-|\alpha_{V}|},\quad B-s_{V}A\simeq\frac{2}{\pi^{2}}\left(H+\frac{2s_{V}Q_{V}}{s}\right)e^{-|\alpha_{V}|},\quad\sigma^{\omega}_{A}-s_{V}\sigma^{\omega}_{H}\simeq\frac{2T^{2}}{\pi^{2}}e^{-|\alpha_{V}|}.$ (33b)
In order to find the angular frequencies supported by Eq. (30), we must solve the equation ${\rm det}\,\mathbb{M}=0$. Writing
$\frac{1}{T^{2}}\mathbb{M}=\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\frac{i}{3\tau_{A}}\mathbb{I}_{A}+\frac{i}{3\tau_{H}}\mathbb{I}_{H},$ (34)
with $\mathbb{I}^{A}_{\ell\ell^{\prime}}=\delta_{\ell A}\delta_{\ell^{\prime}A}$ and $\mathbb{I}^{H}_{\ell\ell^{\prime}}=\delta_{\ell H}\delta_{\ell^{\prime}H}$, one can obtain ${\rm det}\left(\frac{1}{T^{2}}\mathbb{M}\right)$ as
${\rm det}\left(\frac{1}{T^{2}}\mathbb{M}\right)={\rm det}(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega})+\frac{i}{3\tau_{A}}\left|\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}B\\ -\kappa_{\Omega}B&\frac{2\omega}{T^{2}}\sigma_{A}^{\omega}\end{array}\right|\\ +\frac{i}{3\tau_{H}}\left|\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}A\\ -\kappa_{\Omega}A&\frac{2\omega}{T^{2}}\sigma_{A}^{\omega}\end{array}\right|-\frac{2\omega}{9T^{2}\tau_{A}\tau_{H}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)=0\,.$ (35)
@@
${\rm det}(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega})=\left(\frac{2\omega}{T^{2}}\right)^{3}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\\ -\frac{2\omega\kappa_{\Omega}^{2}}{HT^{2}}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}].$ (36)
@@
${\rm det}\left(\omega\mathbb{M}_{v}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\frac{i}{3\tau_{H}}\mathbb{I}_{H}+\frac{i}{3\tau_{A}}\mathbb{I}_{A}\right)\\ =\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)\left(\frac{1}{\tau_{H}}+\frac{1}{\tau_{A}}\right)-\frac{\kappa_{\Omega}^{2}}{H}\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)\right]\\ -\frac{2\omega}{9T^{2}\tau_{A}\tau_{H}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)=0.$ (37)

### v2

@@ 2.3 Linear perturbations
Following Ref. Morales-Tejera:2024uzg , we take small perturbations on top of the rigidly-rotating state. For definiteness, we consider only longitudinal perturbations along the vorticity vector. The detailed derivation of the equations for the linear perturbations is presented in Ref. Morales-Tejera:2024uzg , and for the sake of brevity, we do not repeat it here. Instead, we list the results:
@@
where $D=u^{\mu}\partial_{\mu}$ represents the comoving derivative, $\nabla^{\mu}=\Delta^{\mu\nu}\partial_{\nu}$ is the spatial gradient in the fluid rest frame, while $\theta=\partial_{\mu}u^{\mu}$ is the expansion scalar. As we already discussed in Sec. 2.2 at the level of the charge currents, all three currents $J^{\mu}_{V/A/H}$ would be conserved in a quantum field theory of free (non-interacting) massless fermions. This property is certainly true also in the quantum case for the vector current when $\partial_{\mu}J^{\mu}_{V}=0$. Contrary to Ref. Morales-Tejera:2024uzg , in this paper, we take into account the non-conservation of the axial and helical charges in the so-called relaxation-time approximation.
@@
where $f$ represents a background (possibly vanishing) constant value, and $\omega(k)$ represent the angular frequencies determined by the system of Eqs. (10). [^3] It is important not to confuse the angular frequency $\omega=\omega(k)$, which determines the time evolution of fluctuations (11), with the vorticity $\omega^{\mu}$ which enters, for example, the system of equations (10) being associated with the local angular velocity of the vortical fluid. In this paper, we take the approach of Ref. Morales-Tejera:2024uzg and consider at leading order only longitudinal perturbations. We note that, due to the vortical background, the amplitudes of the transverse velocity perturbations become non-vanshing at first-order with respect to $\Omega$.
In order for linear perturbations to be applicable, we consider an unpolarized background state in which $\mu_{A}=\mu_{H}=0$. Then, the right-hand sides of Eqs. (10d) and (10e) are also linear in perturbations. The presence of the axial and helical relaxation times do not affect the conservation of energy and momentum. Therefore, the equations $\partial_{\mu}T^{\mu\nu}=0$ lead to the usual acoustic modes, characterized by the dispersion relation
@@
with $c_{s}$ being the speed of sound in an ultrarelativistic ideal fluid. Eq. (12) is valid to $O(\Omega^{2})$, as pointed out in Ref. Morales-Tejera:2024uzg , and it is consistent with the results derived in Ref. Abbasi:2016rds . Our result also coincides with that derived in Ref. Gorbar:2017toh , as we are considering the case of a vanishing background axial chemical potential. Note however that our result remains the same even at finite axial chemical potential, while that of Ref. Gorbar:2017toh exhibits a correction proportional to $\Omega Q_{A}$, due to two factors: 1) at finite $Q_{A}$, our reference frame is boosted with a velocity $V=-Q_{A}\Omega/4P$ with respect to the laboratory frame considered in Ref. Gorbar:2017toh ; 2) The amplitudes $\delta u^{x}_{\omega}$ and $\delta u^{y}_{\omega}$ of the transverse components of the velocity pick up an $(x,y)$ dependence at first order with respect to $\Omega$, which was neglected in Ref. Gorbar:2017toh , and which affects the dispersion relation.
The perturbations in the energy-momentum sector are not influenced by perturbations in the charges, $\delta Q_{\ell}$. For the charge-related modes, $\delta P_{\omega}=\delta u_{\omega}=0$ (see discussion in Ref. Morales-Tejera:2024uzg ), such that the equations for the charge currents become:
$\bigl{(}\omega\delta Q_{\ell;\omega}-k\Omega\delta\sigma^{\omega}_{\ell;\omega}\bigr{)}{\biggl{\rvert}}_{\delta P_{\omega}=0}=-\frac{iT^{2}\delta\mu_{\ell;\omega}}{3\tau_{\ell}}.$ (13)
The above equation involves the variations $\delta Q_{\ell;\omega}$ and $\delta\sigma^{\omega}_{\ell;\omega}$ of the charge densities and vortical conductivities at constant pressure. Thinking in terms of the grand canonical ensemble parameters, the constraint $\delta P_{\omega}=0$, together with the thermodynamic relation $\delta P_{\omega}=s\delta T_{\omega}+\sum_{\ell}Q_{\ell}\delta\mu_{\ell;\omega}$ lead to the constraint
$\delta T_{\omega}=-\sum_{\ell}\frac{Q_{\ell}}{s}\delta\mu_{\ell;\omega}.$ (14)
@@
$\mathbb{M}_{\ell\ell^{\prime}}\delta\mu_{\ell^{\prime};\omega}=0,$
$\frac{1}{T^{2}}\mathbb{M}=\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\frac{i}{3\tau_{A}}\mathbb{I}_{A}+\frac{i}{3\tau_{H}}\mathbb{I}_{H},$ (15a)
where we introduced the parameter
$\kappa_{\Omega}=\frac{k\Omega}{T}\,,$ (15b)
which has a sense of a normalized momentum. We also introduced the following notations:
$\mathbb{M}^{\omega}_{\ell\ell^{\prime}}$ $=\frac{1}{T^{2}}\left(\frac{\partial Q_{\ell}}{\partial\mu_{\ell^{\prime}}}-\frac{3Q_{\ell}Q_{\ell^{\prime}}}{sT}+\frac{Q_{\ell^{\prime}}\vec{\mu}}{sT}\cdot\frac{\partial Q_{\ell}}{\partial\vec{\mu}}\right),$ (15c)
$\mathbb{M}^{\Omega}_{\ell\ell^{\prime}}$ $=\frac{1}{T}\left(\frac{\partial\sigma^{\omega}_{\ell}}{\partial\mu_{\ell^{\prime}}}-\frac{2\sigma^{\omega}_{\ell}Q_{\ell^{\prime}}}{sT}+\frac{Q_{\ell^{\prime}}\vec{\mu}}{sT}\cdot\frac{\partial\sigma^{\omega}_{\ell}}{\partial\vec{\mu}}\right),$ (15d)
as well as $\mathbb{I}^{A}_{\ell\ell^{\prime}}=\delta_{\ell A}\delta_{\ell^{\prime}A}$ and $\mathbb{I}^{H}_{\ell\ell^{\prime}}=\delta_{\ell H}\delta_{\ell^{\prime}H}$.
@@ 2.4 Unpolarized background
The pressure corresponding to an ensemble of non-interacting, massless fermions and anti-fermions with polarizations $\lambda=\pm 1/2$ is given by Ambrus:2019khr ; Morales-Tejera:2024uzg :
@@
$q^{V}_{\sigma,\lambda}=\sigma,\qquad q^{A}_{\sigma,\lambda}=2\lambda,\qquad q^{H}_{\sigma,\lambda}=2\lambda\sigma.$ (18)
@@
$\sigma^{\omega}_{\ell;\beta}$ $=\frac{1}{2}\frac{\partial Q_{\ell}}{\partial\mu_{A}}=-\frac{T^{2}}{2\pi^{2}}\sum_{\sigma,\lambda}2\lambda q_{\sigma,\lambda}^{\ell}{\rm Li}_{2}(-e^{\mu_{\sigma,\lambda}/T}).$ (20)
@@
${\rm Li}_{2}(-e^{\alpha})+{\rm Li}_{2}(-e^{-\alpha})$ $=-\frac{\pi^{2}}{6}-\frac{\alpha^{2}}{2},$ (23)
@@
$P$ $=\frac{7\pi^{2}T^{4}}{180}+\frac{\mu_{V}^{2}T^{2}}{6}+\frac{\mu_{V}^{4}}{12\pi^{2}},$
$Q_{V}$ $=\frac{\mu_{V}T^{2}}{3}+\frac{\mu_{V}^{3}}{3\pi^{2}},\ \ \ \sigma_{A}^{\omega}=\frac{T^{2}}{6}+\frac{\mu_{V}^{2}}{2\pi^{2}},$
$\sigma_{H}^{\omega}$ $=\frac{T^{2}}{\pi^{2}}\left[{\rm Li}_{2}(-e^{-\mu_{V}/T})-{\rm Li}_{2}(-e^{\mu_{V}/T})\right],$ (24)
@@
Taking into account the following expressions for the derivatives of the charge densities and vortical conductivities with respect to the chemical potentials,
$\frac{\partial Q_{\ell}}{\partial\mu_{\ell^{\prime}}}{\biggl{\rvert}}_{\mu_{A}=\mu_{H}=0}=2\begin{pmatrix}\sigma^{\omega}_{A}&0&0\\ 0&\sigma^{\omega}_{A}&\sigma^{\omega}_{H}\\ 0&\sigma^{\omega}_{H}&\sigma^{\omega}_{A}\end{pmatrix},$
$\frac{\partial\sigma^{\omega}_{\ell;\beta}}{\partial\mu_{\ell^{\prime}}}=\frac{1}{\pi^{2}}\begin{pmatrix}0&\mu_{V}&TL\\ \mu_{V}&0&0\\ TL&0&0\end{pmatrix},$ (26)
with $L=2\ln\left(2\cosh\frac{\mu_{V}}{2T}\right)$, we find for $\mathbb{M}_{\omega}$ and $\mathbb{M}_{\Omega}$ the following expressions Morales-Tejera:2024uzg :
$\mathbb{M}_{\omega}$ $=\frac{2}{T^{2}}\begin{pmatrix}\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H&0&0\\ 0&\sigma^{\omega}_{A}&\sigma^{\omega}_{H}\\ 0&\sigma^{\omega}_{H}&\sigma^{\omega}_{A}\end{pmatrix},$
$\mathbb{M}_{\Omega}$ $=\begin{pmatrix}0&\frac{1}{H}A&\frac{1}{H}B\\ A&0&0\\ B&0&0\end{pmatrix}.$ (27)
In the above, $\Delta H=H-1$ with $H=(E+P)/(sT)=1+\mu_{V}Q_{V}/(sT)$, while $A$ and $B$ are defined as
@@
$\sigma_{A}^{\omega}=\frac{T^{2}}{6}+\frac{\mu_{V}^{2}}{2\pi^{2}},\qquad\sigma^{\omega}_{H}=\frac{2\ln 2}{\pi^{2}}T^{2}\alpha_{V}+O(\alpha_{V}^{3}),$
$H=1+\frac{15\alpha_{V}^{2}}{7\pi^{2}}+O(\alpha_{V}^{4}),$
$L=2\ln 2+O(\alpha_{V}^{2}),\quad A=\frac{2\alpha_{V}}{7\pi^{2}}+O(\alpha_{V}^{3}),$
$B=\frac{2\ln 2}{\pi^{2}}+\frac{7\pi^{2}-120\ln 2}{28\pi^{4}}\alpha_{V}^{2}+O(\alpha_{V}^{4}).$ (29)
At large chemical potential, when $\lvert\alpha_{V}\rvert\rightarrow\infty$, we have
$H=\frac{\alpha_{V}^{2}}{\pi^{2}}\left(1+\frac{23\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right),$
$A=\frac{2\alpha_{V}}{3\pi^{2}}\left(1-\frac{4\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right),$
$\frac{4s_{V}Q_{v}}{\pi^{2}s}=\frac{4s_{V}}{\pi^{2}\alpha_{V}}(H-1)\simeq$
$\frac{4s_{V}\alpha_{V}}{\pi^{4}}\left(1+\frac{8\pi^{2}}{15\alpha_{V}^{2}}+O(\alpha_{V}^{-4})\right),$ (30a)
where $s_{V}={\rm sgn}(\mu_{V})$ of $\mu_{V}$ is the sign of the vector chemical potential. In what concerns the quantities $L$, $B$ and $\sigma_{H}$, their large-$\alpha_{V}$ behaviour receives exponentially-damped corrections of the form $e^{-\lvert\alpha_{V}\rvert}$, as shown below:
$L\simeq\lvert\alpha_{V}\rvert+2e^{-\lvert\alpha_{V}\rvert},$
$B-s_{V}A\simeq\frac{2}{\pi^{2}}\left(H+\frac{2s_{V}Q_{V}}{s}\right)e^{-\lvert\alpha_{V}\rvert},$
$\sigma^{\omega}_{A}-s_{V}\sigma^{\omega}_{H}\simeq\frac{2T^{2}}{\pi^{2}}e^{-\lvert\alpha_{V}\rvert}.$ (30b)
We are now ready to solve ${\rm det}\mathbb{M}=0$. Starting from Eq. (15a), we can compute the determinant of $\mathbb{M}$ by taking the following steps:
${\rm det}\left(\frac{1}{T^{2}}\mathbb{M}\right)={\rm det}(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega})\\ +\frac{i}{3\tau_{A}}\left\lvert\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}B\\ -\kappa_{\Omega}B&\frac{2\omega}{T^{2}}\sigma_{A}^{\omega}\end{array}\right\rvert\\ +\frac{i}{3\tau_{H}}\left\lvert\begin{array}[]{cc}\frac{2\omega}{T^{2}}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)&-\frac{\kappa_{\Omega}}{H}A\\ -\kappa_{\Omega}A&\frac{2\omega}{T^{2}}\sigma_{A}^{\omega}\end{array}\right\rvert\\ -\frac{2\omega}{9T^{2}\tau_{A}\tau_{H}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)=0\,.$ (31)
@@
${\rm det}(\omega\mathbb{M}_{\omega}-\kappa_{\Omega}\mathbb{M}_{\Omega})=\\ \left(\frac{2\omega}{T^{2}}\right)^{3}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\\ -\frac{2\omega\kappa_{\Omega}^{2}}{HT^{2}}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}].$ (32)
@@
${\rm det}\left(\omega\mathbb{M}_{v}-\kappa_{\Omega}\mathbb{M}_{\Omega}+\frac{i}{3\tau_{H}}\mathbb{I}_{H}+\frac{i}{3\tau_{A}}\mathbb{I}_{A}\right)\\ =\frac{2\omega}{T^{2}}\left\{\left(\frac{2\omega}{T^{2}}\right)^{2}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)[(\sigma^{\omega}_{A})^{2}-(\sigma^{\omega}_{H})^{2}]\right.\\ \left.-\frac{\kappa_{\Omega}^{2}}{H}[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma^{\omega}_{H}]\right\}\\ +\frac{i}{3}\left[\left(\frac{2\omega}{T^{2}}\right)^{2}\sigma_{A}^{\omega}\left(\sigma_{A}^{\omega}-\frac{T^{2}}{3}\Delta H\right)\left(\frac{1}{\tau_{H}}+\frac{1}{\tau_{A}}\right)\right.\\ \left.-\frac{\kappa_{\Omega}^{2}}{H}\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)\right]\\ -\frac{2\omega}{9T^{2}\tau_{A}\tau_{H}}\left(\sigma^{\omega}_{A}-\frac{T^{2}}{3}\Delta H\right)=0.$ (33)